# Emergence of the Primacy Effect in Structured State-Space Models

### v4

@@ 3.2 Models
The models used for the binary memory verification task comprised three layers, as illustrated in Fig. 2. In the first layer, the input integers were embedded into 256-dimensional real-valued vectors. These embeddings were shared between study items and query tokens. The resulting sequence of vectors was then processed by the SSM/RNN, whose outputs were linearly projected onto binary logits to determine whether each query token was present in the study items.
@@
This study primarily examined the single-layer S4 model as the goldstandard implementation of the SSM (Gu et al., 2022b, ).666 Recent studies have shown that the state matrix ($A$) of S4 can be simplified into a purely diagonal form without compromising performance (S4D; Gu et al., 2022a, ). By contrast, the original S4 model introduced an additional low-rank component to the diagonal structure (referred to as the Diagonal Plus Low Rank form, or DPLR) to ensure a mathematically well-founded state matrix. Notably, the diagonal variant exhibited a qualitatively similar primacy effect to the DPLR model. Due to the page limitations, results for the diagonal model are omitted from this paper, and all reported findings are based on the DPLR model. The model encoded the channel-wise dynamics of the input embeddings in a complex-valued space, with its outputs subsequently projected back into the real domain by discarding imaginary components. The state and input matrices ($A$ and $B$ in Eq. 1) were initialized to approximate each channel’s trajectory using Legendre/Laguerre polynomials of degrees 0–63 (HiPPO-LegS/LagT) or a Fourier basis $\{s_{0},c_{0},\dots,s_{31},c_{31}\}$, where $s_{n}(t):=\sqrt{2}\sin(2\pi nt)$ and $c_{n}(t):=\sqrt{2}\cos(2\pi nt)$ (HiPPO-Fout, Fourier Recurrent Unit; Zhang et al.,, 2018; Gu et al.,, 2020, 2023). The matrices were discretized by the bilinear method (Tustin,, 1947).
@@
For comparison, a single-layer long short-term memory (LSTM) network was also evaluated (Hochreiter and Schmidhuber,, 1997). LSTM has been the goldstandard RNN architecture for various time-series processing tasks, including language modeling (Sundermeyer et al.,, 2012; Graves,, 2013). The dimensionality of both hidden and cell states was set to 256.
Another well-established baseline worth considering is the Transformer architecture (Vaswani et al.,, 2017), equipped with a sufficiently long memory window. However, it turned out that the Transformer can easily achieve perfect accuracy on the adopted task—even at the maximal levels of input length and vocabulary size implementable within the available computational resources—thereby failing to incur the memory load necessary for evaluating the primacy effect777 For the same reason, the proposed experimental paradigm was also inadequate for testing extended architectures of the SSM (Mamba; Gu and Dao,, 2024; Dao and Gu,, 2024) and the LSTM (xLSTM; Beck et al.,, 2024), both of which attained perfect accuracy at the highest task difficulty settings. Consequently, the Transformer is excluded from the main results reported in the next section; instead, the Appendix provides a detailed explanation of how it can solve the memory verification task, with particular emphasis on the role of its attention mechanism.
@@
The models were trained for 300,000 iterations using the Adam optimizer with parameters $(\beta_{0},\beta_{1}):=(0.9,0.99)$ (Kingma and Ba,, 2015). Batch size was set to 512. The learning rate was linearly increased from 0.0 to 0.001 over the first 1,000 iterations (warmups) and subsequently decayed according to the cosine annealing schedule Loshchilov and Hutter, (2017). To prevent gradient explosion, the gradient norm was clipped at 1.0. The Python code for the experiments is available at https://github.com/An0nym0usAuth0r/NeuralPrimacyEffect.git.
@@ 4.1 Emergence of the Primacy Effect
Fig. 3 reports the accuracy of the binary memory verification task across all combinations of memorization and verification times. The brightness of each cell in the square heatmaps indicates the accuracy for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. That is, they report the proportion of true positives against false negatives (i.e., the recall score). Additionally, the top separate row of each panel displays the accuracy for distractor queries (integers not included among the study items), capturing the prevalence of true negatives over false positives. Just below it, the second row summarizes the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across all verification times (columns).
The binary memory verification performance of the SSM model was highest for study items presented at the beginning of the sequence, demonstrating a clear primacy effect (Fig. 33). The model maintained high accuracy across different query timings (as indicated by the bright colors in the top rows of the heatmaps), provided that the sequence length did not exceed its capacity (see the accuracy decline in Fig. 3, where $L=256$). In other words, memory for the initial study items exhibited minimal decay over time.
By contrast, the LSTM did not display this primacy effect; its accuracy was uniform across both the memorization and verification phases (Fig. 3).
These findings held true regardless of whether the state and input matrices of the SSM were optimized for the task (Fig. 3) or remained fixed at their initial values (Fig. 33). Moreover, the results remained consistent across different polynomial bases underlying the state and input matrices, including Laguerre (Fig. 3), Fourier (3), and Legendre (all other panels).
@@
Figure 3: Accuracy of the binary memory verification task. Each cell in the square heatmaps represents the accuracy (or the recall score) for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the SSM (S4) model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ (i.e., $0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
Figure 4: Top row: Optimization trajectories of the discretization step size $\Delta t$ in the SSM (S4) model with frozen Legendre state and input matrices. Each heatmap column represents the distribution of $\Delta t$ across 256 latent channels $\times$ 10 training runs, sorted in the ascending order. Bottom row: Histograms displaying the initial (blue) and final (orange) values of $\Delta t$, aggregated across the 256 latent channels $\times$ 10 training runs (see Fig. 5 in the Appendix for variations among individual runs). The first to third panel columns from the left show the results for three different ranges of log-uniformly random initializations, whereas the fourth and fifth columns tested shorter and longer study items, respectively.
@@ 4.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ plays a critical role in determining the memory capacity of the SSM. Moreover, once the state matrix $A$ is fixed, $\Delta t$ becomes the sole parameter capable of influencing the dynamics of the SSM;888 Freezing $\Delta t$ resulted in a complete failure of learning. all remaining parameters are confined to feedforward transforms.
Accordingly, to further investigate its role, the optimization trajectories of $\Delta t$ were tracked over the course of training. The analysis revealed that as training progressed, a specific range of step sizes ($\Delta t\leq 0.03$) became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig. 4).
@@ 5 Discussions
5.1 Summary of the Findings
The present study demonstrated that the SSM exhibits the primacy effect in memorization. When performing the binary memory verification task, which parallels paradigms used to investigate memory capacity in humans and animals (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985), the model showed the highest accuracy for study items presented at the beginning of the sequence. Moreover, memorized information was not retrievable immediately after the presentation of the study items. These findings are novel and counterintuitive, as they challenge the theoretical formulation of the SSM, which assumes an exponentially decaying measure (for Legendre and Laguerre bases; Fig 1; Gu et al.,, 2023).
As noted in the Preliminaries, the SSM was designed to achieve a longer-lasting memory than classical RNNs (Gu et al.,, 2020). Prior research on RNNs and SSMs has focused on their ability to preserve input data against temporal decay. A typical benchmark for these inquiries is the delayed reconstruction task (a.k.a. the copying memory task; Arjovsky et al.,, 2016). In this task, models are provided with a short sequence of integers—usually ten repeatable random samples ranging between 0 and 7—followed by null inputs (e.g., zero vectors), and are then asked to reconstruct the initially observed sequence in the original order. Model performance is evaluated based on the length of the null-valued interval between the initial observations and their reconstructions.
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit.999 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$, as all the other trainable parameters pertain exclusively to feedforward transforms. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\Delta t\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
@@ 5.1 Summary of the Findings
5.2 Implications for Computational Neuroscience
Most existing theories of the primacy effect in human and animal memory have focused on abstract, procedural explanations of the cognitive bias. One influential account is the dual-store hypothesis (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968), which attributes the primacy and recency effects to two distinct memory systems. According to this view,study items presented early in a sequence are transferred from short-term memory—where all incoming observations are temporally buffered—into long-term memory, where they are preserved against decay, thereby producing the primacy effect. By contrast, the recency effect is thought to arise from the residual information retained in short-term memory, which has limited capacity and only holds the most recent inputs.
Moving beyond such abstract accounts, several researchers have proposed biologically grounded neural network models that replicate the primacy effect (Burgess et al.,, 1991; Wong et al.,, 1991; Greene et al.,, 2000; Sikström,, 2006; Lansner et al.,, 2013). The present findings align with this line of work but offer a novel perspective. Specifically, the SSM examined here was developed with the aim of advancing practically useful artificial intelligence, independently of any deliberate attempt to model the primacy effect from a psychological/biological standpoint. Thus, the spontaneous emergence of the primacy effect in the SSM presents a unique and complementary hypothesis for the mechanism underlying this cognitive bias, opening new avenues for future investigation in computational neuroscience.
Importantly, the industrial orientation of the SSM does not imply biological implausibility. Voelker and Eliasmith, (2018) and Voelker et al., (2019) established a connection between the SSM and spiking neural networks, demonstrating that the dynamics of the state vector in the SSM (Eq. 1) can be implemented using a population of spiking neurons, which remain the most plausible model of biological neuronal activity to date. Moreover, Voelker and Eliasmith, (2018) showed that the spiking instantiation of the SSM, when tuned to implement some particular function known as the delay line, reproduces the characteristic behavior of biological time cells,101010 Time cells fire at specific moments following onset cues/events and are believed to encode temporal information (MacDonald et al.,, 2011). which have been observed in several brain regions, including the hippocampus (MacDonald et al.,, 2011; Eichenbaum,, 2014), striatum (Mello et al.,, 2015), and cortex (Luczak et al.,, 2015). Comparing a spiking implementation of the SSM exhibiting the primacy effect with brain recordings from psychological experiments may thus yield novel insights into the neuro-computational mechanisms underlying this cognitive phenomenon.
Of course, the present study has certain limitations as a biological simulation. Most notably, the model here was explicitly optimized for the memorization task from scratch, whereas the primacy effect observed in humans and animals is presumably an outcome of their natural cognitive development. Thus, fundamentally different mechanisms may underlie the similar memorization biases exhibited by the SSM and biological organisms. Future studies may address this discrepancy by training the model in a more biologically naturalistic setting. However, as discussed in the previous section, training an ANN on empirical data, such as natural language, introduces the risk of conflating human-originated cognitive biases—transferred through the data—with the model’s intrinsic properties.
It is also important to note that the SSM does not constitute a comprehensive model of sequence memorization; its lack of the recency effect suggests the necessity of incorporating an additional module to account for short-term memory processes. Accordingly, future work should investigate whether the SSM can reliably function as a long-term memory component when integrated with a complementary short-term memory mechanism (e.g., finite-queue/Markov systems), either through modular combination or ensemble learning.
The model’s capacity also far exceeded that of biological memory systems. The binary memory verification task employed in this study involved extremely long lists of study items ($L\in\{64,128,256\}$) as well as a huge vocabulary ($K=4096$), which is unlikely to be replicated in human or animal experiments. This level of task difficulty was necessary to clearly demonstrate the primacy effect in the SSM, and was still insufficient to reach the memory limits of more powerful architectures such as Mamba (Gu and Dao,, 2024; Dao and Gu,, 2024) and xLSTM (Beck et al.,, 2024),thereby constraining the scope of model comparison here. Future research should explore downscaling the model toward more biologically plausible conditions.
Finally, the present study does not offer an explanation for various well-documented characteristics of the psychological primacy effect. For instance, the retention of initial study items is thought to be achieved, or at least facilitated, by rehearsal—repetitive mental or verbal recitation of observed items before requested recall/verification (Rundus,, 1971; Marshall and Werder,, 1972; Glenberg et al.,, 1980)—for which the models studied here lack an analogous mechanism.
In addition, the models operated on a unitless discrete time scale, to which the continuous-time representation of the SSM was aligned. As a result, they are unsuitable for investigating physical-time effects, such as the impact of retention intervals between study-item presentation and verification (Cornell and Bergstrom,, 1983; Wright et al.,, 1985; Neath,, 1993) or variations in inter-item intervals (Neath and Knoedler,, 1994). This limitation could be addressed by exploring spiking neural networks equivalent the SSM. As noted above, the SSM can be converted into a spiking neural network, which in turn enables the simulation of the physical-time effects as well as the incorporation of biologically inspired connectivity structures (Voelker and Eliasmith,, 2018).
@@ Acknowledgments
This study was supported by JST ACT-X (JPMJAX21AN) and Core Research for Evolutional Science and Technology (JPMJCR22P5); JSPS Grant-in-Aid for Early-Career Scientists (JP21K17805) and for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the ACCMS, Kyoto University, regarding the use of their supercomputer system.
Appendix A Appendix
@@
A.1 Remarks on Transformer
As noted in the Methods section, the present study was unable to assess whether or not Transformer exhibits the primacy effect (without inheriting human-induced biases). Specifically, the model consistently achieved perfect accuracy on the binary memory verification task, thereby failing to encounter the memory load necessary for evaluating the primacy effect. This section offers a theoretical account of how Transformer is able to solve this task, with particular focus on the role of its attention mechanism.
to begin, recall that the “memory” in Transformer (and other finite-queue/Markov models in general) consists of the raw input observation, preserved without decay or lossy compression. Although this approach is highly non-economic for processing long sequences, growing computational resources have made it viable in practice.
Given full access to the uncompressed sequence of prior observations, Transformer computes attention weights based on the dot-product similarity between (linear projected) input embeddings. In the context of the memory verification task, this computation enables a direct comparison between the study items and verification queries; setting aside the interference from positional encodings, the resulting attention distribution at each verification time step becomes sharply peaked (one-hot-like) when the query matches a study item, or almost uniform when it does not. As a result, Transformer can trivially solve the task by exploiting this attention distribution.
A.2 Distributions of $\Delta t$ across Training Runs
Fig. 5 reports the distribution of the $\Delta t$ parameters before and after each of the ten training runs (decomposing the leftmost panel in Fig. 4).

### v5

@@ 3.2 Models
The models used for the binary memory verification task comprised three layers, as illustrated in Figure 3. In the first layer, the input integers were embedded into 256-dimensional real-valued vectors. These embeddings were shared between study items and query tokens. The resulting sequence of vectors was then processed by the SSM, whose outputs were linearly projected onto binary logits to determine whether each query token was present in the study items.
@@
This study primarily examined the single-layer S4 model as the goldstandard implementation of the SSM (Gu et al., 2022b, ).777 Recent studies have shown that the state matrix ($A$) of S4 can be simplified into a purely diagonal form without compromising performance (S4D; Gu et al., 2022a, ). By contrast, the original S4 model introduced an additional low-rank component to the diagonal structure (referred to as the Diagonal Plus Low Rank form, or DPLR) to ensure a mathematically well-founded state matrix. Notably, the diagonal variant exhibited a qualitatively similar primacy effect to the DPLR model. Due to the page limitations, results for the diagonal model are omitted from this paper, and all reported findings are based on the DPLR model. The model encoded the channel-wise dynamics of the input embeddings in a complex-valued space, with its outputs subsequently projected back into the real domain by discarding imaginary components. The state and input matrices ($A$ and $B$ in Eq. 3) were initialized to approximate each channel’s trajectory using Legendre/Laguerre polynomials of degrees 0–63 (HiPPO-LegS/LagT) or a Fourier basis $\{s_{0},c_{0},\dots,s_{31},c_{31}\}$, where $s_{n}(t):=\sqrt{2}\sin(2\pi nt)$ and $c_{n}(t):=\sqrt{2}\cos(2\pi nt)$ (HiPPO-Fout, Fourier Recurrent Unit; Zhang et al.,, 2018; Gu et al.,, 2020, 2023). The matrices were discretized by the bilinear method (Tustin,, 1947).
@@
For comparison, a single-layer long short-term memory (LSTM) network was also evaluated (Hochreiter and Schmidhuber,, 1997). The dimensionality of both hidden and cell states was set to 256.
@@
The models were trained for 300,000 iterations using the Adam optimizer with parameters $(\beta_{0},\beta_{1}):=(0.9,0.99)$ (Kingma and Ba,, 2015). Batch size was set to 512. The learning rate was linearly increased from 0.0 to 0.001 over the first 1,000 iterations (warmups) and subsequently decayed according to the cosine annealing schedule (Loshchilov and Hutter,, 2017). To prevent gradient explosion, the gradient norm was clipped at 1.0. The Python code for the experiments is available at https://github.com/tkc-morita/primacy-effect.git.
@@ 4.1 Emergence of the Primacy Effect
Figure 4: Accuracy of the binary memory verification task. Each cell in the square heatmaps represents the accuracy (or the recall score) for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (4) depicts the accuracy distribution for the LSTM, while the other panels (44) report results for the SSM (S4) under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 4 and 4, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 4, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 4, where the upper bound was extended to $1.0$ (i.e., $0.001\leq\Delta t\leq 1.0$). The length of study items was set to $L=128$, except in panel 4 ($L=64$) and panel 4 ($L=256$).
Figure 4 reports the accuracy of the binary memory verification task across all combinations of memorization and verification times. The brightness of each cell in the square heatmaps indicates the accuracy for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. That is, they report the proportion of true positives against false negatives (i.e., the recall score). Additionally, the top separate row of each panel displays the accuracy for distractor queries (integers not included among the study items), capturing the prevalence of true negatives over false positives. Just below it, the second row summarizes the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across all verification times (columns).
The binary memory verification performance of the SSM model was highest for study items presented at the beginning of the sequence, demonstrating a clear primacy effect (Figure 44). The model maintained high accuracy across different query timings (as indicated by the bright colors in the top rows of the heatmaps), provided that the sequence length did not exceed its capacity (see the accuracy decline in Figure 4, where $L=256$). In other words, memory for the initial study items exhibited minimal decay over time.
By contrast, the LSTM did not display this primacy effect; its accuracy was uniform across both the memorization and verification phases (Figure 4).
These findings held true regardless of whether the state and input matrices of the SSM ($A$ and $B$ in Eq. 3) were optimized for the task (Figure 4) or remained fixed at their initial values (Figure 44). Moreover, the results remained consistent across different polynomial bases underlying the state and input matrices, including Laguerre (Figure 4), Fourier (4), and Legendre (all other panels).
@@ 4.2 Distribution of the Time-Step Sizes
Figure 5: Top row: Optimization trajectories of the discretization step size $\Delta t$ in the SSM (S4) with frozen Legendre state and input matrices. Each heatmap column represents the distribution of $\Delta t$ across 256 latent channels $\times$ 10 training runs, sorted in the ascending order. Bottom row: Histograms displaying the initial (blue) and final (orange) values of $\Delta t$, aggregated across the 256 latent channels $\times$ 10 training runs (see Figure 6 in Appendix A for variations among individual runs). The first to third panel columns from the left show the results for three different ranges of log-uniformly random initializations, whereas the fourth and fifth columns tested shorter and longer study items, respectively.
As discussed in §2.1.2, the discretization time-step size $\Delta t$ plays a critical role in determining the memory capacity of the SSM. Moreover, once the state matrix $A$ is fixed, $\Delta t$ becomes the sole parameter capable of influencing the dynamics of the SSM;888 Freezing $\Delta t$ resulted in a complete failure of learning. all remaining parameters are confined to feedforward transforms.
Accordingly, to further investigate its role, the optimization trajectories of $\Delta t$ were tracked over the course of training. The analysis revealed that as training progressed, a specific range of step sizes ($\Delta t\leq 0.03$) became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Figure 5).
@@ 5 Discussions
The present study demonstrated that the SSM exhibits the primacy effect in memorization. When performing the binary memory verification task, which parallels paradigms used to investigate memory capacity in humans and animals, the model showed the highest accuracy for study items presented at the beginning of the sequence. Moreover, memorized information was not retrievable immediately after the presentation of the study items. These findings are novel and counterintuitive, as they challenge the theoretical formulation of the SSM, which assumes an exponentially decaying measure (for Legendre and Laguerre bases; Fig 2; Gu et al.,, 2020, 2023).
As noted in §2.1.1, the SSM was designed to achieve a longer-lasting memory than classical RNNs (Gu et al.,, 2020). Prior research on RNNs and SSMs has focused on their ability to preserve input data against temporal decay. However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit.999 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$, as all the other trainable parameters pertain exclusively to feedforward transforms. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\Delta t\leq 0.03$). As discussed in §2.1.2, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains open why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory. Future research may address this issue through comparisons across a wider range of discretization methods—such as the Runge-Kutta method—extending beyond the standard empirical options of bilinear and zero-order hold.
@@
It also remains an open question whether the primacy effect holds in more advanced settings than those examined in this study. Specifically, the scope was restricted to single-layered models, whereas empirical applications almost invariably employ multi-layered architectures. Such extended architectures achieved perfect accuracy on the adopted task—even at the maximal levels of input length and vocabulary size implementable within the available computational resources—thereby failing to incur the memory load necessary for evaluating the primacy effect. For the same reason, the proposed experimental paradigm was also inadequate for testing the SSM-based language model, Mamba (Gu and Dao,, 2024; Dao and Gu,, 2024). Therefore, Clarifying whether the primacy effect persists in such powerful architectures is an important avenue for future research.
@@ Acknowledgments
This study was supported by JST AIP Accelerated Program (JPMJCR25U6), ACT-X (JPMJAX21AN), and Core Research for Evolutional Science and Technology (JPMJCR22P5); JSPS Grant-in-Aid for Early-Career Scientists (JP21K17805) and for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.
Appendix A Distributions of $\Delta t$ across Training Runs
Figure 6 reports the distribution of the $\Delta t$ parameters before and after each of the ten training runs (decomposing the leftmost panel in Figure 5).
@@ Appendix A Distributions of $\Delta t$ across Training Runs
Appendix B Associative Recall
This section investigates the primacy effect in a more advanced task—associative recall (Fu et al.,, 2023)—compared to the binary memory verification. Similar to the memory verification setup, the associative recall task first presents a sequence of random tokens (study items) to the model, followed by a shuffled version of the same sequence (excluding the final token; Figure 7).101010 In the original formulation of the associative recall task, the model is presented with a single query token sampled from the study items. This study extends the task in a more empirical setting, requiring the model to recall multiple pairs of precedent and successor tokens within a single sequence. Unlike the verification task, however, none of the shuffled query tokens are replaced with distractors; instead, the model is required to return the immediate successor of each query token. For example, given study items $(8,29,2,17)$ and queries $(2,8,29)$, the correct outputs are $(17,29,2)$.
[…]
Figure 7: Schematic illustration of the (extended) associative recall task.
Despite its apparent simplicity, previous work has shown that the ability to recall the successor of previously presented input tokens is a meaningful indicator of language models’ capacity (Olsson et al.,, 2022; Fu et al.,, 2023; Gu and Dao,, 2024). As such, investigating associative recall helps bridge the primacy effect observed in SSMs with real-world applications, particularly in language modeling.
Model and task hyperparameters were manually tuned to ensure that overall accuracy remained suboptimal. Specifically, the length of the study items was set to $L=96$, and the vocabulary size to $K=128$. The number of latent channels in the SSM (also equal to the dimensionality of the input embeddings) was set to 1024. All other hyperparameters were identical to those used in the memory verification task (see §3.1).
[…]
Figure 8: Accuracy of the associative recall task performed by the SSM (S4) under different parameter configurations. Each cell in the square heatmaps represents the accuracy (or the recall score) for study items (target successors) that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The top separate row and the right separate column display the average accuracy across memorization times (rows) and recall times (columns), respectively. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using the Legendre (8,8), Laguerre (4), and Fourier (4) polynomials. The state and input matrices were optimized for the task in panel 4, and remained fixed at their initial values elsewhere. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$. The length of study items was set to $L=96$, and the vocabulary size to $K=128$.
Figure 8 shows the SSM’s accuracy on the associative recall task. The model exhibited higher accuracy for initial study items compared to terminal ones (vertical differences), despite overall retention being less persistent than in the memory verification task (horizontal declines). Besides this general trend, the model also appears to have learned delayed reconstructions of input sequences, as evidenced by the diagonal patterns observed in the heatmaps. This behavior may reflect a statistical property of the adopted task; input-output distances of length $L$ occurred more frequently than either shorter or longer dependencies during training (see Footnote 6).